OpenStudy (anonymous):

Show that lim g(x)=f'(x), f'(x) = -1/x^2, g(x) = (1/(x+h)-1/x)/h

hartnn (hartnn):

did u try this ?where r u stuck ?

OpenStudy (anonymous):

Umm I'm not really sure where to begin.

hartnn (hartnn):

ok, so u need to find \(\large \lim \limits_{x \rightarrow 0} \dfrac{\dfrac{1}{x+h}-\dfrac{1}{x}}{h}\) right ?

hartnn (hartnn):

can u solve numerator part ?

OpenStudy (anonymous):

yes. to show that it is equal to \[\frac{ -1 }{x ^{2} }\]

OpenStudy (anonymous):

1/x-1/x=0? so h/h, which = 1?

hartnn (hartnn):

sorry, it should be \(\large \lim \limits_{h \rightarrow 0} \dfrac{\dfrac{1}{x+h}-\dfrac{1}{x}}{h}\) and no, how u got that ?

hartnn (hartnn):